OpenStudy (anonymous):

If g(x) = 2(x − 4), find the value of x if g(x) = 20. 32 12 14 10

mathslover (mathslover):

\[\huge\color{blue}{2(16)=32}\]

OpenStudy (unklerhaukus):

\[g(x) = 2(x − 4) = 20. \]\[\qquad(x-4)=10\]\[\qquad \qquad x=14\]